OpenStudy (rainbow_dash):

Find the value of cos(x), if sec(x) = 3/2. 0

OpenStudy (anonymous):

secant is the inverse of cosine. Likewise, cosine is the inverse of secant. Therefore, \[\sec(x) = 3/2\] \[cosx = (secx)^{-1} = \frac{ 2 }{ 3 }\]

OpenStudy (rainbow_dash):

thanks :)